# Leveraging Atom Loss Errors in Fault Tolerant Quantum Algorithms

### v2

@@ 3 Techniques for addressing qubit loss in deep circuits
To effectively handle loss errors in deep circuits prior to logical measurement, we now explore several SE methods that simultaneously manage loss and Pauli errors. These methods differ in their qubit overhead, gate operations, and loss detection capabilities. We focus on three primary approaches: (1) conventional SE, which resembles traditional circuit-based quantum computing (CBQC) with various modifications for delayed-erasure conversion, (2) teleportation-based SE, which resembles measurement-based quantum computing (MBQC) [54, 55] and (3) modified Steane SE [32]. We find that the underlying circuits and resulting performance of the methods are similar, but not identical: all approaches successfully remove qubit loss and have trade-offs in logical error rates and qubit overheads depending on the specific ratios between loss and Pauli errors. We begin by detailing each SE approach before analyzing numerical results in Section 3.4. In Section 3.5, we relate these performance differences to key metrics such as lifecycle length and number of entangling gates to predict performance based on the noise model for each SE method.
@@ 3.1 Modified conventional SE
The conventional SE method involves repeated stabilizer measurements using physical measure qubits. Data qubits are not directly measured or replaced when lost, and thus over time, logical performance can degrade in the presence of loss. We consider augmenting conventional SE by utilizing physical SWAP gates to detect losses on all qubits (SWAP SE), as proposed in Refs. [56, 42] and further explored in Refs. [24, 23]. At the end of each SE round, a SWAP gate and physical SWAP movement are performed between data and measure qubits. Conveniently, this approach does not require any additional entangling gates by applying gate cancellation identities (see Fig. 3(a)). If the data qubit is not lost, the SWAP operations cancel each other, such that the resulting measure qubit stores the stabilizer outcome. Conversely, if the data qubit was lost, the loss is directly identified through the SSR measurement and automatically replaced. Thus, using this method, each qubit cycles through both roles of data and measure qubits, ensuring a uniform lifecycle length of $\sim 8$ for all qubits in the bulk, even in deep circuits (see Supplementary Materials  [50] Section S2.1). However, a data qubit loss detection comes at the cost of losing a stabilizer outcome information, potentially impacting performance, as observed numerically in the next section.
Each SWAP operation incurs a cost due to qubit movement, making it crucial to evaluate different SWAP periods to balance performance and experimental complexity. The SWAP period defines how often loss-detecting SE rounds are interspersed with conventional SE rounds, occurring at a fixed interval. We identify a trade-off between the lifecycle lengths of data and measurement qubits for different SWAP periods, demonstrating that a period of 2 can be competitive with a period of 1 (see Supplementary Materials [50] Figs. S2, S3). A period of 2 reduces experimental complexity while maintaining a similar average lifecycle length.
@@ 3.2 Teleportation-based SE
Teleportation-based SE is an alternative QEC approach which utilizes regular teleportation between logical qubit layers, inherently enabling loss detection upon logical measurement [54, 55, 22, 16] (see Fig. 3(b)). While typically associated with measurement-based quantum computation (MBQC) architectures, the teleportation-based SE can be used in a circuit-based framework where gates are directly applied on qubits. We realize the XZZX cluster state [30], as described in detail in the Supplementary Materials [50] Section S2.5. In this framework, each qubit is entangled with four neighboring qubits, which, under our error model considering noise on entangling gates only, results in short lifecycles of length four. Among the SSR-based approaches considered, this SE method enables the shortest lifecycles with no experimental overhead apart from SSR, though it requires additional qubits. These features make it well-suited for neutral atom architectures and particularly effective in loss-dominant regimes, though its advantages diminish when the loss fraction is small, as detailed in Section 3.4.
3.3 Modified Steane SE
Steane SE uses transversal $CX$ gates applied to fault-tolerant logical measure qubits to extract syndromes and detect errors [32]. Transversal gates, which apply the same operation across corresponding physical qubits in each code block, ensure that errors propagate predictably from logical data qubits to logical measure qubits. If the measure qubits are prepared fault-tolerantly with $d$ SE rounds, one round of Steane SE suffices to accurately capture the syndromes in a given basis [32]. However, conventional Steane SE lacks a native mechanism for detecting loss, as losses on data qubits are not detected by the logical measure qubits and remain hidden until the end. We modify the conventional Steane SE approach by incorporating logical SWAP operations (SWAP gate and SWAP movement), as illustrated in Fig. 3(c), thereby enabling loss detection through SSR at each logical measurement. This approach leverages the teleportation of logical information and resembles Knill SE [57].
The modified Steane SE can be thought of as an interpolation between modified conventional SE and teleportation-based SE, similar to the ideas in Ref. [58]. In the limit of multiple logical measure qubits, each prepared with one SE round, the modified Steane SE becomes equivalent to teleportation-based SE (see Supplementary Materials [50] Section S2.6). Conversely, multiple rounds of SE can be used to prepare a higher-quality logical measure qubit, providing flexibility to adapt to various error models and resource constraints. Steane SE also supports pre-selection of measure qubit blocks based on quality, potentially increasing fidelity without significant overhead [59, 60, 61, 62]. To provide a comprehensive perspective on SE techniques, we include Steane SE here despite not numerically analyzing it in the main text. Since modified Steane SE interpolates between modified conventional SE and teleportation-based SE, we expect its performance to qualitatively reflect this interpolation. Thus, we highlight its conceptual advantages as a potential direction for future studies.
Figure 3: Loss-detecting SE methods. (a) Modified SWAP SE method for detecting and correcting loss errors using physical SWAP operations between data and measure qubits, without gate overhead. (b) Teleportation-based SE method, utilizing layers of teleportation and stabilizer checks in alternating bases. The circuit is repeated $d$ times. (c) Steane SE method with logical SWAP, employing logical fault-tolerant measurement qubits and teleportation. (d) Logical error rates as a function of SE rounds using the delayed-erasure decoder, with a code distance of 7 and a physical error rate $1\%$. Comparable performance is observed across SE methods in regimes with short lifecycles. (e) Error thresholds as a function of the loss fraction for different SE methods, showcasing the improvement with loss for all methods. (f) Effective distance as a function of loss fraction for different SE methods, considering distance $d=7$. (g) Space-time overhead as a function of the loss fraction, describing the required distance to achieve logical error rates of $10^{-12}$ given a physical error rate of $0.5\%$. The legend for (d-g) appears in (e).
@@ 3.4 Comparison of SE methods
We now compare the performance of these approaches in the context of a surface code memory, using circuit-level simulations with varying loss fractions $L={p_{\text{loss}}}/({p_{\text{loss}}+p_{\text{Pauli}}})$, where $p_{\text{loss}}$ and $p_{\text{Pauli}}$ are physical loss and Pauli error probabilities, respectively. We identify the optimal SE approach at different loss fractions and Pauli error biases (see Supplementary Materials [50] Section S4 for detailed results with bias).
We also evaluate a Free SE method, which detects and replaces lost qubits at different intervals: after every gate (period 0.25) or every $K$ rounds (period $K$). Free detection at period 0.25 eliminates loss propagation, ensuring that the exact loss locations are known, effectively realizing an erasure channel. Furthermore, we compare our results to a Free SE method with period 1, including an additional feature: the exact loss locations are known, greatly simplifying the decoding problem. A possible implementation of these SE approaches involves mid-circuit erasure conversion and atom replacement [17].
@@
Figure 3(d) compares the logical error rates of conventional SE without loss detection, approaches with loss detection—such as SWAP SE, teleportation-based SE, and Free SE—and Free SE with perfect loss information, in a single logical memory experiment at an experimentally-motivated loss fraction of 0.5. For a small number of SE rounds ($\leq 10$), all protocols behave similarly. However, as the number of SE rounds increases, loss detection and the replacement of lost qubits substantially enhance performance, providing a convenient approach to achieving high circuit depths without additional experimental gate overhead.
@@
Figure 3(e) shows the threshold of each SE method as a function of the loss fraction. We determine the threshold by calculating the logical error rate for different code distances $d$, using a noiseless initialization, followed by $d-1$ rounds of noisy stabilizer measurements and a final noiseless transversal measurement. We find that the thresholds of all methods improve with increasing loss fraction. Notably, methods with shorter lifecycles, such as teleportation-based SE, benefit more from increasing loss fraction. In the next Section 3.5, we provide further insight into the thresholds of each SE method by linking it to simple characterizations such as qubit lifecycles.
@@
Figure 3(f) presents the effective code distance of each SE method as a function of the loss fraction, for distance $d=7$. The effective distance $d_{e}$ refers to the number of errors required to cause a logical failure, as the logical error rate scales as $(p/p_{\text{th}})^{d_{e}}$ far below the threshold $p_{\text{th}}$, where $p$ is the physical error rate. For Pauli noise, $d_{e}=(d+1)/2$, and for erasure noise $d_{e}=d$. To determine $d_{e}$ for each loss fraction, we fit the logical error rate data far below the threshold to the function $\alpha p^{\beta}$, where $\alpha$ and $\beta$ are fitting parameters (see Supplementary Materials [50] Fig. S13). As expected, all SE methods, while utilizing the delayed-erasure decoder, experience increased effective distance with an increased loss fraction. For loss errors only, teleportation-based SE and Free SE methods (periods 0.25 and 1) achieve the optimal effective distance of $d_{e}\approx d$, while SWAP SE achieves $d_{e}\approx d-1$, likely due to longer lifecycles. Additional threshold and effective distance results under an alternative error model from Ref. [22] are presented in Table 1.
@@
To account for the qubit overhead required in different SE methods, we evaluate the space-time overhead for performing $d$ SE rounds of each method in Figure 3(g). For a given physical error rate of $0.5\%$, we determine the required code distance to achieve a logical error rate of $10^{-12}$ by fitting the logical error as a function of distance for $d=3,5,7,9$. We then present the space-time overhead of each approach as a function of the loss fraction, highlighting the overhead reduction as the loss fraction increases for all SE approaches. Our results highlight that although teleportation-based SE requires more qubits, its better error suppression when the loss fraction is high leads to a more favorable space-time volume compared to SWAP SE (see space-time overhead for each SE method in Table 1).
An important regime is where the loss fraction is $\approx 0.5$, which is comparable to the ratio observed in recent neutral atom experiments [14, 10]. In this regime, the thresholds of the different methods range from 1.5% to 2.5%, and all SE methods achieve similar effective distances. However, these methods can vary significantly in experimental complexity. In particular, erasure conversion requires directly detecting and replacing the qubit erasure without projectively measuring it, unlike with SSR [17]. Clearly, all methods provide benefits when the loss is correctly managed and detected, and experimental considerations can illuminate which may work best in practice.
Another important aspect of neutral atom qubits is their intrinsic bias, as $Z$-type Pauli errors are much more common than $X$-type Pauli errors [14]. In Supplementary Materials Section S4.2.2 [50], we investigate the interplay between biased errors and delayed erasure errors across multiple SE methods, employing the XZZX surface code [28] and the XZZX cluster state [30] for teleportation-based SE. Our analysis considers scenarios both with and without bias-preserving gates, as well as the presence of biased erasure noise. Bias-preserving gates, such as native $CX$ gates, preserve the bias of the errors, in contrast to $CX$ gates decomposed into $CZ$ and $H$ gates. Biased erasure noise occurs when a qubit exits the computational subspace exclusively from the state |$\ket{1}$. As a result, the replaced qubit follows a biased error channel instead of a depolarizing error channel. Additionally, we investigate scenarios where Pauli errors are biased, examining the two-dimensional space defined by the bias ratio and the loss fraction. Thresholds are calculated at each point in this space for various SE methods (see Figures S16 and S17 in the Supplementary Materials [50]). A key observation emerges across all SE methods: increasing the loss fraction has a significantly greater impact on thresholds than increasing the bias ratio, even when using bias-preserving gates and biased-erasure. This difference in impact between the loss fraction and the bias ratio becomes even more pronounced in the absence of bias-preserving gates.
@@ 3.5 Predicting performance by error counting
The curve intersections with the axes provide key insights. The y-axis intersection, representing the threshold in the absence of Pauli errors, is related to the lifecycle length. For the Free SE period 1 with perfect loss information, the threshold is $p_{\text{loss,th}}\approx 7.2\%$. Increasing the lifecycle reduces the threshold, which decays with the lifecycle length, as shown in Fig. 4(b) (see Supplementary Materials [50] Section S2.7 for further details). Notably, the SWAP SE, with an average lifecycle length of 8, has a lower threshold of $p_{\text{loss,th}}\approx 2.3\%$, deviating from the heuristic. This lower performance is attributed to additional factors unique to SWAP SE, such as the inability to detect measurement qubit loss when data qubits are detected as lost and the loss of stabilizer information, as discussed previously in Section 3.1.
The x-axis intersection, representing Pauli error thresholds, correlates with the number of entangling gates, which impacts the number of Pauli errors in the final state. For example, over $d$ SE rounds, teleportation-based SE uses 1.5×1.5$\times$1.5 × more entangling gates than SWAP SE and Free SE approaches, resulting in a proportional decrease in the Pauli error threshold, as similarly observed in Ref. [16].
@@ 4 The effect of Loss Errors in deep Logical Algorithms design
Figure 5: Deep logical circuits with qubit loss. (a) A deep Clifford logical algorithm consisting of random logical single-qubit and transversal $CX$ gate layers, with periodic SE rounds at varying frequencies. (b, c) Circuit-level simulation results showing the logical error rate as a function of the number of SE rounds per transversal gate layer, for different loss fractions. ($p=1\%$, $d=5$, 24 layers). The SWAP SE method (c) effectively mitigates loss errors, restoring the optimal SE frequency observed in Pauli-dominated scenarios. In contrast, conventional SE (b) exhibits varying error correction regimes, where loss can either improve or degrade performance depending on the lifecycle length.
Loss errors in logical algorithms can have significantly different effects compared to standard memory benchmarks. We now study the effects of loss errors on the QEC design of multi-qubit deep logical algorithms. Specifically, we analyze how physical loss errors influence the optimal frequency of SE rounds between transversal gates. Our analysis reveals that SWAP-based SE achieves comparable optimal SE frequencies in both the presence and absence of loss.
We focus on the effects of loss in random Clifford logical algorithms (Fig. 5(a)), where multiple logical qubits interact through transversal gates interspersed with periodic SE rounds at tunable frequencies. Circuit-level simulations are performed for logical circuits comprising 24 layers of logical $CX$ and single-qubit logical gates ($X,Y,Z$), with SE rounds applied at varying intervals, as in Ref. [12]. Two scenarios are analyzed: conventional SE without explicit loss detection, which detects only measurement qubit loss in each SE round (Fig. 5(b)), and SWAP-based SE (Fig. 5(c)). The x-axis represents the number of SE rounds per gate layer, ranging from no SE ($n_{r}=1/24$) to multiple SE rounds after every gate layer ($n_{r}=3$).

### v3

@@ 3 Techniques for addressing qubit loss in deep circuits
In deep logical circuits prior to logical measurement, qubit lifecycles are extended, introducing many correlated errors and degrading performance. While the delayed-erasure decoder mitigates these effects, its effectiveness deteriorates when lifecycles become too long due to error accumulation. To maintain high performance, loss detection and qubit replacement must be integrated into the QEC process. In this section, we explore practical syndrome extraction (SE) methods that enable such replenishment with minimal overhead. We primarily focus on: (i) conventional SE, which resembles traditional circuit-based quantum computing (CBQC) with various modifications for delayed-erasure conversion utilizing SSR or erasure conversion capabilities, and (ii) teleportation-based SE, which resembles measurement-based quantum computing (MBQC) [58, 59]. Additionally, a description of a modified Steane SE approach, which bridges conventional and teleportation-based SE but is not simulated here, is provided in Appendix Section D.D.4. These methods differ in their qubit overhead, gate operations, loss detection capabilities, and experimental requirements, as analyzed below.
We find that the underlying circuits and resulting performance of the methods are similar, but not identical: all approaches successfully remove qubit loss and have trade-offs in logical error rates and qubit overheads depending on the specific ratios between loss and Pauli errors. We begin by detailing each SE approach before analyzing numerical results in Section 3.4. In Section 3.5, we relate these performance differences to key metrics such as lifecycle length and number of entangling gates to predict performance based on the noise model for each SE method.
@@ 3.1 Modified conventional SE
The conventional SE method involves repeated stabilizer checks using physical measure qubits. Data qubits are not directly measured or replaced when lost, and thus over time, logical performance can degrade in the presence of loss. We consider augmenting conventional SE by utilizing SSR and physical SWAP gates to detect losses on all qubits (SWAP SE), as proposed in Refs. [60, 42] and further explored in Refs. [25, 24]. At the end of each SE round, a SWAP gate and physical SWAP movement are performed between data and measure qubits. Conveniently, this approach does not require any additional entangling gates by applying gate cancellation identities (see Fig. 3(a)). If the data qubit is not lost, the SWAP operations cancel each other, such that the resulting measure qubit stores the stabilizer outcome. Conversely, if the data qubit was lost, the loss is directly identified through the SSR measurement and automatically replaced. Thus, using this method, we correct losses of all types of qubits, each cycles through both roles of data and measure qubits, ensuring a uniform lifecycle length of $\sim 8$ for all qubits in the bulk even in deep circuits (see Appendix D.1). However, a data qubit loss detection comes at the cost of losing a stabilizer outcome information, potentially impacting performance, as observed numerically in the next section.
Each SWAP operation incurs a cost due to qubit movement, making it crucial to evaluate different SWAP periods to balance performance and experimental complexity. The SWAP period defines how often loss-detecting SE rounds are interspersed with conventional SE rounds, occurring at a fixed interval. We identify a trade-off between the lifecycle lengths of data and measurement qubits for different SWAP periods, demonstrating that a period of 2 can be competitive with a period of 1 (see Figures. 1415). A period of 2 reduces experimental complexity while maintaining a similar average lifecycle length.
@@ 3.2 Teleportation-based SE
Teleportation-based SE uses repeated logical teleportation to fresh code blocks, inherently detecting loss via SSR at each step [58, 59, 23, 17] (see Fig. 3(b)). In our implementation, logical qubits are prepared in one initialization round and then entangled in the XZZX pattern, which is equivalent to the XZZX cluster state used in MBQC [31] (See Appendix D.D.3 for more details). While typically associated with MBQC, it can also be implemented in a circuit-based framework where gates are directly applied to qubits. Under our error model with noise only on entangling gates, this construction yields short lifecycles of length four for all physical qubits in the bulk. Among the SSR-based approaches considered, it achieves the shortest lifecycles without experimental operations beyond SSR, though at the cost of additional qubits. These features make it well-suited for neutral-atom architectures and particularly effective in loss-dominant regimes, as discussed in Sec. 3.4.
3.3 Mid-circuit erasure conversion SE approaches
Mid-circuit erasure conversion offers an alternative to SSR-based detection by converting physical qubit loss into flagged erasure events during circuit execution (See Fig. 3(c).) This has been demonstrated in alkaline-earth-like atoms [18, 11, 17, 12] and superconducting qubits [20, 22, 21, 16, 37, 38, 39, 50, 61], using auxiliary fluorescence imaging, auxiliary qubits, or leakage monitors, followed by reset or transport. The resulting erasure is often delayed and requires active intervention, such as qubit replacement or reset, imposing experimental overhead in real-time implementations.
We refer to this as the direct conversion SE method, and explore multiple detection-replacement schedules: (i) detection and replacement after each gate layer or SE round (periods 0.25,1), and (ii) detection after each gate but replacement only at the end of the round (period 1 + loss moment information). Frequent detection improves decoder performance; frequent replacement reduces logical error by halting loss propagation, but might extend circuit duration due to feed-forward control. This flexibility is a practical advantage of erasure conversion.
Figure 3: Loss-detecting SE methods. (a) Modified SWAP SE: data and measure qubits are swapped each round to detect loss via SSR; the SWAP’s CNOT decomposition cancels with existing parity-extraction gates, with the remaining CNOT replaced by classical feedforward, so no additional entangling gates are required. (b) Teleportation-based SE: logical qubits are teleported to fresh blocks prepared in alternating bases, detecting loss via SSR and shortening lifecycles, at the cost of extra qubits. (c) Direct conversion SE: mid-circuit measurement and replacement convert loss to erasure using additional hardware [18]; detection/replacement frequency is varied in simulations. (d) Logical error rates vs. SE rounds ($d=7$, $p=1\%$) using the delayed-erasure decoder. Comparable performance is observed across SE methods in regimes with short lifecycles. (e) Error thresholds vs. loss fraction, showcasing the improvement with loss for all SE methods. (f) Effective distance vs. loss fraction ($d=7$), improving with loss for all SE methods. (g) Space-time overhead to reach $P_{L}=10^{-12}$ at $p=0.5\%$ vs. loss fraction. Legends for (d-g) appear in (e).
@@ 3.4 Comparison of SE methods
We now compare the performance of these approaches in the context of a surface code memory, using circuit-level simulations with varying loss fractions $L={p_{\text{loss}}}/({p_{\text{loss}}+p_{\text{Pauli}}})$, where $p_{\text{loss}}$ and $p_{\text{Pauli}}$ are physical loss and Pauli error probabilities, respectively. We identify the optimal SE approach at different loss fractions and Pauli error biases (see Appendix Section C for detailed results with bias).
@@
Figure 3(d) compares the logical error rates of conventional SE without loss detection, approaches with loss detection—such as SWAP SE, teleportation-based SE, and direct conversion SE—and direct conversion SE with perfect loss information, in a single logical memory experiment at an experimentally-motivated loss fraction of 0.5, comparable to that observed in recent neutral atom experiments [15, 11, 62, 10]. For a small number of SE rounds ($\leq 10$), all protocols behave similarly. However, as the number of SE rounds increases, loss detection and the replacement of lost qubits substantially enhance performance, providing a convenient approach to achieving high circuit depths without additional experimental gate overhead.
@@
Figure 3(e) shows the threshold of each SE method as a function of the loss fraction. We determine the threshold by calculating the logical error rate for different code distances $d$, using a noiseless initialization, followed by d1d-1 rounds of noisy stabilizer checks and a final noiseless transversal measurement. We find that the thresholds of all methods improve with increasing loss fraction. Notably, methods with shorter lifecycles, such as teleportation-based SE, benefit more from increasing loss fraction. In the next Section 3.5, we provide further insight into the thresholds of each SE method by linking it to simple characterizations such as qubit lifecycles.
@@
Figure 3(f) presents the effective code distance of each SE method as a function of the loss fraction, for distance $d=7$. The effective distance $d_{e}$ refers to the number of errors required to cause a logical failure, as the logical error rate scales as $(p/p_{\text{th}})^{d_{e}}$ far below the threshold $p_{\text{th}}$, where $p$ is the physical error rate. For Pauli noise, $d_{e}=(d+1)/2$, and for erasure noise $d_{e}=d$. To determine $d_{e}$ for each loss fraction, we fit the logical error rate data far below the threshold to the function $\alpha p^{\beta}$, where $\alpha$ and $\beta$ are fitting parameters (see Fig. 24). All SE methods, while utilizing the delayed-erasure decoder, experience increased effective distance with an increased loss fraction. For loss errors only, teleportation-based SE and direct conversion SE methods (periods 0.25 and 1) achieve the optimal effective distance of $d_{e}\approx d$, while SWAP SE achieves $d_{e}\approx d-1$, likely due to longer lifecycles. Additional threshold and effective distance results under an alternative error model from Ref. [23] are presented in Table 1.
@@
To account for the qubit overhead required in different SE methods, we evaluate the space-time overhead for performing $d$ SE rounds of each method in Figure 3(g). We define space-time overhead as the product of the total number of physical qubits and the circuit depth required for the computation. This metric provides a platform-independent measure of logical performance. However, it does not include experimental considerations such as real-time qubit transport, additional laser resources, or mid-circuit measurement overhead. These factors can significantly affect implementation and are discussed qualitatively in Sections 3.1-3.3. For a given physical error rate of $0.5\%$, we determine the required code distance to achieve a logical error rate of $10^{-12}$ by fitting the logical error as a function of distance for $d=3,5,7,9$. We then present the space-time overhead of each approach as a function of the loss fraction, highlighting the overhead reduction as the loss fraction increases for all SE approaches. Our results highlight that although teleportation-based SE requires more qubits, its better error suppression when the loss fraction is high leads to a more favorable space-time volume compared to SWAP SE (see space-time overhead for each SE method in Table 1).
At the experimentally relevant loss fraction of $\sim 0.5$, seen in multiple different hardware systems such as neutral atoms and ions  [15, 11, 62, 10], the thresholds of the different methods range from 1.5% to 2.5%, and all SE methods achieve similar effective distances. However, their experimental feasibility differs. Direct conversion SE requires erasure conversion with mid-circuit detection and real-time replacement of lost qubits, often demanding extra lasers and fast feedback. By contrast, SSR-based approaches such as SWAP or teleportation SE avoid these requirements. Thus, while all methods benefit from accurate loss detection, the optimal choice depends strongly on hardware-specific capabilities and constraints.
[…]
@@ 3.5 Predicting performance by error counting
The curve intersections with the axes provide key insights. The y-axis intersection, representing the threshold in the absence of Pauli errors, is related to the lifecycle length. For the direct conversion SE period 1 with perfect loss information, the threshold is $p_{\text{loss,th}}\approx 7.2\%$. Increasing the lifecycle reduces the threshold, which decays with the lifecycle length, as shown in Fig. 4(b). Notably, the SWAP SE, with an average lifecycle length of 8, has a lower threshold of $p_{\text{loss,th}}\approx 2.3\%$, deviating from the heuristic. This lower performance is attributed to additional factors unique to SWAP SE, such as the loss of stabilizer information when the swapped in data qubit is lost and the fact that measurement qubit loss is only detected in the following round, after it propagates as a data qubit loss, as discussed in Section 3.1.
The x-axis intersection, representing Pauli error thresholds, correlates with the number of entangling gates, which impacts the number of Pauli errors in the final state. For example, over $d$ SE rounds, teleportation-based SE uses $1.5\times$ more entangling gates than SWAP SE and direct conversion SE approaches, resulting in a proportional decrease in the Pauli error threshold, as similarly observed in Ref. [17].
@@ 4 The effect of Loss Errors in deep Logical Algorithms design
Figure 5: Deep logical circuits with qubit loss. (a) A deep Clifford logical algorithm consisting of random logical single-qubit and transversal CXCX gate layers, with periodic SE rounds at varying frequencies. (b, c) Circuit-level simulation results showing the logical error rate as a function of the number of SE rounds per transversal gate layer, for different loss fractions. ($p=1\%$, $d=5$, 24 layers). The SWAP SE method (c) effectively mitigates loss errors, restoring the optimal SE frequency observed in Pauli-dominated scenarios. In contrast, conventional SE (b) exhibits varying error correction regimes, where loss can either improve or degrade performance depending on the lifecycle length.
Loss errors in logical algorithms can have significantly different effects compared to standard memory benchmarks. We now study the effects of loss errors on the QEC design of multi-qubit deep logical algorithms. Specifically, we analyze how physical loss errors detection influence the optimal frequency of SE rounds between transversal gates. Our analysis reveals that SWAP-based SE achieves comparable optimal SE frequencies in both the presence and absence of loss.
We focus on the effects of loss in random Clifford logical algorithms (Fig. 5(a)), where multiple logical qubits interact through transversal gates interspersed with periodic SE rounds at tunable frequencies. Circuit-level simulations are performed for logical circuits comprising 24 layers of logical $CX$ and single-qubit logical gates ($X,Y,Z$), with SE rounds applied at varying intervals, as in Ref. [13] (for more details see Appendix I). Two scenarios are analyzed: conventional SE without explicit loss detection, which detects only measurement qubit loss in each SE round (Fig. 5(b)), and SWAP-based SE (Fig. 5(c)). The x-axis represents the number of SE rounds per gate layer, ranging from no SE ($n_{r}=1/24$) to multiple SE rounds after every gate layer ($n_{r}=3$).